Integrate dx/(4+2²)½ dx
1
Expert's answer
2020-02-18T06:38:31-0500

"\frac{x}{(4+2^2)^{1/2}}dx"  


"=\frac{1}{(4+2^2)^{1/2}} \int xdx"

"=\frac{x^2}{2} \cdot \frac{1}{(4+2^2)^{1/2}}+C"

"=\frac{x^2}{2}\cdot \frac{1}{2 \cdot 2^{1/2}}+C"

"=\frac{x^2}{4 \cdot 2^{1/2}}+C"

Answer: "\frac{x^2}{4 \cdot 2^{1/2}}+C"